OpenStudy (anonymous):

Let the function f be defined by f(x)=x^2 + 18. If m is a positive number such that f(2m)=2f(m), what is the value of m?

jimthompson5910 (jim_thompson5910):

f(x)=x^2 + 18 f(2m)=(2m)^2 + 18 f(2m) = 4m^2 + 18

jimthompson5910 (jim_thompson5910):

f(x)=x^2 + 18 2*f(m)=2*(m^2 + 18) 2*f(m)=2m^2 + 36

jimthompson5910 (jim_thompson5910):

f(2m)=2f(m) 4m^2 + 18 = 2m^2 + 36 keep going to solve for m

OpenStudy (anonymous):

haha thank you! that really helped explain how to do it very well!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

M would equal 3?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

okay I think I understand how to do this now.... thank you again!

jimthompson5910 (jim_thompson5910):

np